THE MILLIMETER/SUBMILLIMETER SPECTRUM OF CYANOPHOSPHINE, PH_{2}CN (X̃^{1}A′)

\underline{D. T. HALFEN}, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ 85721; D. J. CLOUTHIER, Department of Chemistry, University of Kentucky, Lexington, KY 40506; and L. M. ZIURYS, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ 85721.

Cyanophosphine, PH_{2}CN (X̃^{1}A′), is the third row analog of cyanamide, a known interstellar molecule, and could be a potential interstellar species. However, only the microwave spectrum has been measured to date, and predictions at millimeter/submillimeter wavelengths have high uncertainities. Therefore, the millimeter/submillimeter spectrum of cyanophosphine has been measured using direct absorption techniques. This molecule was created by the reaction of gas-phase phosphorus, (CN)_{2}, and H_{2} in the presence of argon carrier gas and an AC glow discharge. Eleven transitions have been recorded in the range 305-422 GHz each with multiple asymmetry components from K_{a} = 0 to 8. No evidence of inversion splitting was visible in the spectrum. The data have been fit with an asymmetric top Hamiltonian, and the spectroscopic constants have been determined. The A rotational constant and several higher-order centrifugal distortion parameters have been determined for the first time, allowing accurate predictions in the millimeter/submillimeter region. Here we report our laboratory measurements and an astronomical search.